in an evacuated 2.0 liter container, 1.8 moles of N2 and 2.2 moles of H2 are placed. at equilibrium 0.6 M of NH3 are present. solve for Kc
1
Expert's answer
2020-11-09T14:06:11-0500

N2 + 3 H2 ------------- > 2 NH3

Initial = 1.8 mole 2.2 mole


Eqillibrium = ( 1.8 - x ) mole ( 2.2 - 3x ) mole 2x mole .


Concentration of NH3 = 2x / 2 = 0.6 M

Concentration of N2 = [ N2 ] = 1.8 - x / 2 = 1.8 - 0.6 / 2 = 0.6 M

Concentration of H2 = [ H2 ] = 2.2 - 3 x / 2 = (2.2 - 3 * 0.6 ) / 2 = 0. 2 M


Kc = [ NH3 ]2 / [ H2 ]3 . [ N2 ]


Kc = ( 0.6 )2 / ( 0.6 ) * ( 0.2 )3 .


Kc = 75 .